OpenStudy (anonymous):

how do I simplify 27 1/3 a mixed number?

terenzreignz (terenzreignz):

Simple algorithm/method to do this... \[\huge a\frac{p}q\]

terenzreignz (terenzreignz):

\[\huge a\frac{p}q=\frac{}{\color{red}q}\] Same denominator...

terenzreignz (terenzreignz):

Numerator times whole number... \[\huge a\frac{p}q=\frac{\color{blue}{a\times q}}{q}\]

OpenStudy (anonymous):



OpenStudy (anonymous):